OpenStudy (anonymous):

9(2 + x) = 10(7 + x) answer as a fraction

OpenStudy (anonymous):

9(2 + x) = 10(7 + x) Factor. 18 + 9x = 70 + 10x Subtract 18 from both sides. 9x = 52 + 10x Subtract 10x from both sides. -x = 52 Divide both sides by -1. x= -52 If you want it as a fraction, just put the -52 over one. -52 / 1

OpenStudy (anonymous):

8(7x - 3) = -6 Answer As A Fraction

OpenStudy (anonymous):

8(7x - 3) = -6 56x - 25 = -6 56x = 19 x = 19 / 56